OpenStudy (anonymous):

cot theta 7/18. Where theta lies in quadrant 3. What is csc?

OpenStudy (anonymous):

use your trig identities , 1+cot^2 θ = csc ^2 θ 1+ cot^2 θ = 4 cot^2 θ = 3 cot θ = √3 since quadrant 2 it is negative -√3 2) 1 + 49/324 = csc^2 θ csc θ = √373 / 18 this is also negative.

OpenStudy (anonymous):

i hope i helped